1 p Question 16 Amass of 8kg is traveling at 4m/s. What constant force is required to slow it to a halt in a period of 2 seconds?
media%2Fe88%2Fe8892742-5eb4-45af-963d-56
Answer #1

16) let the force needed is F

Using second law of motion

F* t = change in momentum

F * 2 = 8 * (0 - 4)

F = 16 N

the force needed is 16 N